OpenStudy (anonymous):

cos^-1 rad2/2 cos^-1 (-rad3/2) PLEASE EXPLAIN

OpenStudy (anonymous):

basically the question you need to ask yourself is: what angle of cos gives me rad2/2

OpenStudy (anonymous):

what angle of cos gives me -rad3/2

OpenStudy (anonymous):

its is helpful to remeber the unitcirlce and the know angles

OpenStudy (anonymous):

But I know what those are i just dont understand question like rad 2/2 is 45 and idk would it just be negative 30 for the -rad3/2

OpenStudy (anonymous):

basically just think of this: if cos(45)=rad2/2, then cos inverse (rad2/2)=45

OpenStudy (anonymous):

But i dont understand why what about for -rad3/2 what would inverse cos = -30

OpenStudy (anonymous):

no it would be 5pi/6

OpenStudy (anonymous):

:[ What exactly am i trying to find

OpenStudy (anonymous):

you are trying to find out where cos is negative and equal to rad3/2

OpenStudy (anonymous):

where is cos negative? in what quadrents?